OpenStudy (anonymous):

what is the amplitude period phase shift of f(x)= -3cos(4x +pi) + 6

OpenStudy (anonymous):

@mathmate

myininaya (myininaya):

Do you know how to graph y=cos(x)?

OpenStudy (anonymous):

not really

OpenStudy (mathmate):

In the standard form of f(x)=a*cos(b(x-h))+k the period is \[\frac{2\pi}{|b|}\] phase shift is h, and amplitude is a. If you would transform the given function into its standard from, you can read off these parameters.

OpenStudy (anonymous):

tbh today is the last day of school and im rushing to finish everything before 11:59

OpenStudy (anonymous):

is there like a really fast way to learn this

OpenStudy (jdoe0001):

\(\large f(x)= -{\color{blue}{ 3}}cos({\color{brown}{ 4}}x +{\color{olive}{ \pi}}) + 6\qquad \begin{array}{llll} {\color{blue}{ amplitude}}\\ \textit{phase shift}=\cfrac{{\color{olive}{ \square }}}{{\color{brown}{ \square }}}\\ \textit{period}=\cfrac{2\pi}{{\color{brown}{ \square }}} \end{array}\)

OpenStudy (mathmate):

:)

OpenStudy (anonymous):

ooh ok now i put them in fractions?

OpenStudy (anonymous):

helloooo

OpenStudy (jdoe0001):

well you'd just use the values given in the function, to get the amplitude, phase shift and perod